OpenStudy (anonymous):

A circular lid has a diameter of 12 inches. What is the exact area of the lid? A. 6π square inches B. 12π square inches C. 24π square inches D. 36π square inches

OpenStudy (jordan123321):

what do you think

OpenStudy (igreen):

\(A = \pi r^2\) The radius is half the diameter. 12 / 2 = ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

12/2 = 6

OpenStudy (igreen):

Yes, that is the radius, plug it in the formula: \(A = \pi(6^2)\) Simplify \(6^2\)

OpenStudy (anonymous):

A≈113.1in² i thinkz

OpenStudy (anonymous):

so Simplify 36

OpenStudy (anonymous):

yes

OpenStudy (igreen):

Yes, so we have \(36\pi\).

OpenStudy (igreen):

That's our answer \(\Huge\ddot\smile\)

OpenStudy (anonymous):

thx